1) A corporation is holding its board election at a properly called annual shareholder meeting.  All 10 million outstanding shares are present and voting.   All nine of the current board members are standing for re‐election.  There are no other candidates.   Carlton obtains 3 million shares voting in favor of his re‐election.  Lou obtains 4 million shares voting in favor of his re‐election. Who has been re‐elected?

a. Carlton and Lou. b. Carlton but not Lou. c. Neither Carlton, nor Lou. d. Lou but not Carlton.

2)BBI has 5 million shares of common stock authorized, 1 million shares of common stock issued, and 800,000 shares of outstanding common stock.   How many shares of treasury stock does BBI have?

a. 200,000. b. 1.8 million. c. 4 million. d. 4.2 million.

3)Assume that Hollander, Inc. and NFC, Inc. consolidate into a new corporation called Legend Foods.  Who is now legally responsible for NFC’s existing contractual obligations?

a. Hollander. b. NFC. c. Hollander’s shareholders. d. None of the above.

4)The certificate of incorporation of Mordor, Inc. provides that shareholders have preemptive rights.   Mordor wishes to issue 100,000 shares of common stock to an employee in exchange for future work for Mordor.  Which of the following statements is most accurate?

a. Mordor must first offer these new shares to existing common shareholders. b. Mordor may not issue shares of stock today for a promise of future services. c.The shareholders must vote to approve this transaction. d. This is permissible.
Answer #1

1) Cumulative voting requirement : Total number of votes = 10,000,000 * 9 = 90,000,000

Minimum number of votes to ensure re election = [1 / (9+1) *90,000,000 ] / +1 = 9,000,0001

If Straight voting is used:

Minimum number of votes = (10,000,000 / 2)  + 1 =5,000,0001

Going by both the method , none of the directors have been re ,elected since they did not secure the minimum votes.

Option , c. Neither Carlton, nor Lou.

2) Option a,200,000 [ 1,000,000 - 800,000]

3) Option , b. NFC

4) Option d. This is permissible. [ Pre emptive right applies only when the company intends to issue stock on the stock exchange].